# Feature Request Analysis and Processing: Tasks, Techniques, and Trends

### v1

@@ Abstract.
Feature requests are proposed by users to request new features or enhancements of existing features of software products, which represent users’ wishes and demands. Satisfying users’ demands can benefit the product from both competitiveness and user satisfaction. Feature requests have seen a rise in interest in the past few years and the amount of research has been growing. However, the diversity in the research topics suggests the need for their collective analysis to identify the challenges and opportunities so as to promote new advances in the future. In this work, following a defined process and a search protocol, we provide a systematic overview of the research area by searching and categorizing relevant studies. We select and analyze 131 primary studies using descriptive statistics and qualitative analysis methods. We classify the studies into different topics and group them from the perspective of requirements engineering activities. We investigate open tools as well as datasets for future research. In addition, we identify several key challenges and opportunities, such as: (1) ensuring the quality of feature requests, (2) improving their specification and validation, and (3) developing high-quality benchmarks for large language model-driven tasks.
@@ 1. Introduction
Due to the considerable value of feature requests to software, the analysis and processing of feature requests in software has been widely studied in practice. Recent years have witnessed a surge research interest in feature requests, such as identifying feature requests from issues reports (Iacob and Harrison, 2013; Carreño and Winbladh, 2013; Deocadez et al., 2017; Dhinakaran et al., 2018), extracting software features from users’ requests (Bakar et al., 2016; Bakiu and Guzman, 2017; Johann et al., 2017; Licorish et al., 2017; Dalpiaz and Parente, 2019), feature requests prioritization (Kegel, 2023; de Lima and Marcacini, 2024; Salleh et al., 2024), and locating features in the source code (Palomba et al., 2017; Panichella et al., 2016; Ciurumelea et al., 2017; Kato and Goto, 2017). These have greatly promoted the application of feature requests in software engineering. Furthermore, to investigate the challenges and opportunities in software change request repositories, Cavalcanti et al. (Cavalcanti et al., 2014) conducted a systematic mapping study based on two dimensions: challenges and opportunities related to change request repositories. Tavakoli et al. (Tavakoli et al., 2018) surveyed the state-of-the-art tools and techniques employed in detecting informative aspects in mobile application reviews. Bakar et al. (Bakar et al., 2015) provided a systematic literature review of the state-of-the-art approaches in feature extractions from natural language requirements for reuse in software product line engineering. Wang et al. (Wang et al., 2019) investigated the implicit feedback and explicit feedback employed for crowdsourced requirements engineering (RE) activities. Most existing secondary studies have focused on narrow aspects of feature requests, such as analyzing app store reviews for feature requests or addressing specific tasks like feature extraction. However, to the best of our knowledge, there is no comprehensive study that provides a holistic view of the entire research landscape on feature requests across different tasks and contexts. Furthermore, the field lacks an integrated discussion of overarching challenges and open research opportunities that span the full lifecycle of feature request management. Additionally, there is a lack of understanding regarding the frequency and location of studies published over the past few years.
@@
In this paper, we conduct a systematic literature review on feature requests to portray the landscape of research practices from a requirements engineering perspective and pinpoint research challenges and opportunities, following the guidelines of (Petersen et al., 2008, 2015). We propose four research questions and define a set of search strings to retrieve the initial relevant literature. Then snowballing is employed to retrieve the whole literature. Finally, a total of 131 literature studies are selected. In particular, the studies are scrutinized for feature requests analysis and processing for the software maintenance. We report on the demographics of the published articles to provide a comprehensive view. Furthermore, we also categorize the literature according to the research topic and provide a detailed analysis of each topic. Figure 1 presents an overview of the study topics in relation to requirements activities, encompassing 12 distinct research topics. We also expand each research topic with an explanation of the research progress, accompanied by a demonstration of the proposed approach. To facilitate future studies, we investigate the public tools and datasets for feature requests. Finally, we pinpoint the research gaps, which can help researchers understand the road ahead in investigating user requests.
@@ 2.2. Related Work
Cavalcanti et al. (Cavalcanti et al., 2014) performed a systematic mapping study in 2014 to characterize the research challenges and opportunities for software change request repositories. The result was a two-taxonomy based on two dimensions: challenges and opportunities related to change request repositories. However, the search strings of this paper mainly involve “change request” and “bug report”, etc., not “feature request”. Thus, some feature request-related research may be missed. Tavakoli et al. (Tavakoli et al., 2018) surveyed the state-of-the-art tools and techniques used to detect valuable information concerning user needs and requirements from mobile application reviews. User reviews of mobile applications may contain bug reports, user experience, and user requirements, which are pivotal information relevant to app development and improvement. Therefore, they surveyed the mining tools that have been used to unleash the value of app reviews for mobile app development. They mainly focus on obtaining useful information from user reviews, which our mapping study pays attention to the analysis and processing of user feature requests, which is downstream of their survey. Wang et al. (Wang et al., 2019) provided a mapping study of literature on crowdsourced user feedback employed for requirements engineering purposes. They provided an overview of the types of user feedback that have been employed for crowdsourced requirements engineering activities. They investigated both implicit feedback and explicit feedback. However, they only mapped the literature to requirements engineering activities without further investigation into the state-of-the-art techniques for each topic. Bakar et al. (Bakar et al., 2015) provided a systematic literature review of the state-of-the-art approaches in feature extractions from natural language requirements for reuse in software product line engineering. Bischoff et al. (Bischoff et al., 2019) created a panoramic view of the current literature to pinpoint gaps and supply insights in the integration of feature models. They sought to understand, characterize, and summarize the literature about the integration of feature models. Recently, Cai et al. (Cai et al., 2025) reviewed existing work on requirements elicitation from user-generated content, categorized studies by data sources, methods, and representations, and proposed a research framework to guide future work in this area.
Although these mapping studies or reviews are closely related to feature requests, either regarding how to obtain feature requests from user feedback or about specific software features, however, there is no systematic mapping study or literature review towards the analysis and processing of user feature requests in software engineering. We fill this gap with a systematic literature review study, where we analyze the research topics on feature requests and state-of-the-art techniques within each topic.
@@ 3.2.1. Search Strategy
To identify the maximum number of relevant literature, we employ a well-defined search strategy that combines both manual searching and snowballing. We manually search papers from the databases and digital libraries listed in Table 1 as the starting point of snowballing, with the following search strings:
(“feature request”) OR (“user request”) OR (“user requirement”) OR (“enhancement report”) OR (“enhancement request”) OR (“just-in-time requirement”)
After removing duplicates, we retrieved a total of 17,381 papers from the above databases and digital libraries.
Table 1. List of Digital Libraries.
Digital libraries URLs
IEEE Xplore http://www.ieee.org/web/publications /xplore/
ACM Digital Library http://portal.acm.org
Science Direct http://www.sciencedirect.com/
Springer Link https://link.springer.com/
Wiley InterScience https://onlinelibrary.wiley.com/
Elsevier http://www.elsevier.com
Google Scholar https://scholar.google.com
DBLP https://dblp.org
arXiv https://arxiv.org
@@ 3.2.2. Inclusion/Exclusion Criteria
To identify the most relevant studies on feature request analysis and processing, we drew upon prior work (Giray et al., 2023; Zakari et al., 2020; Croft et al., 2022; Niu et al., [n. d.]; Hu et al., 2025) and established a set of inclusion (ICs) and exclusion criteria (ECs), summarized in Table 2. These criteria were applied to the titles, abstracts, and keywords of candidate papers to ensure relevance and quality.
@@
Through this rigorous screening process, we refined a comprehensive yet focused set of relevant studies for detailed analysis. Overall, our study pool consists of 99 selected papers.
Table 2. Inclusion and Exclusion Criteria
@@
IC3 The paper is published in a peer-reviewed journal, conference proceedings, or a workshop, with the full-text available, including recent publications available on arXiv.
@@
EC1 The paper is about the application of user requests for non-RE activities, like marketing, recommending apps, etc.
@@
EC3 The paper is a technical report or thesis.
3.3. Snowballing
To ensure a comprehensive literature collection, we performed both forward and backward snowballing on our initial set of papers, following the guidelines outlined by Wohlin (Wohlin, 2014). This iterative process involved examining the references and citations of the selected studies in successive rounds. After each round, we applied the established inclusion and exclusion criteria to screen the newly identified articles. Ultimately, after three rounds of snowballing, we finalized a set of 137 papers for our literature review.
3.4. Quality Assessment
@@
Four authors independently assessed all 137 selected studies using this checklist. Whenever discrepancies arose, the reviewers engaged in discussions to reach an agreement. As a result of this thorough evaluation, six papers were excluded for not meeting the established quality criteria, leaving 131 high-quality studies for detailed data extraction and further analysis.
Table 3. Quality Assessment Checklist for Feature Request Studies
@@
3.5. Data Extraction
@@ 4. Demographics
A total of 131 research papers were included in the final study set, covering publications up to August 2025. As shown in Figure 3, the number of primary studies has increased significantly over time, showing a clear upward trend from 2011 to 2017, with the peak occurring in 2017 (16 studies). Activity remained relatively high through 2020, followed by a decline likely due to shifting research priorities. A notable resurgence occurred in 2024 (11 studies), coinciding with the rise of LLMs, which introduced new research opportunities. Overall, the data reveals cyclical research activity with notable peaks and dips over the years. Figure 4 presents the distribution of the selected papers across journals and conferences, highlighting the most frequently appearing publication venues. A total of 84 papers were published in conferences, 41 in journals, and 6 as preprints. Most studies were published in top-tier software engineering outlets, particularly in the field of requirements engineering. High-frequency venues include International Conference on Requirements Engineering (RE, 15 papers), International Conference on Automated Software Engineering (ASE, 8 papers), International Conference on Software Engineering (ICSE, 7 papers), International Working Conference on Requirement Engineering: Foundation for Software Quality (REFSQ, 7papers), Empirical Software Engineering (ESE, 7 papers), Journal of Systems and Software (JSS, 6 papers), International Conference on Software Maintenance and Evolution (ICSME, 4 papers), Requirements Engineering Journal (RE, 4 papers), and International Conference on Evaluation and Assessment in Software Engineering (EASE, 4 papers). These results confirm that the topic has received strong attention from the software engineering community.
@@ 5. Analyzing and Processing Feature Requests
Throughout the lifecycle of feature requests, twelve research topics have been identified in the literature on feature request analysis and processing, covering the entire process. The first step is identifying feature requests from various sources of user feedback, such as issue reports or chat messages. Once identified, the requests undergo quality review and approval prediction to determine their suitability for further processing. During the processing phase, automated techniques may be used to extract key features, recognize user intent, and summarize request content. Classification and prioritization are also applied to support analysis. Traceability links to other artifacts are established to aid in management. After analysis, the feature request is assigned to developers, where techniques like code localization and recommendation help accelerate the implementation process.

### v2

@@ Abstract.
Feature requests are proposed by users to request new features or enhancements of existing features of software products, which represent users’ wishes and demands. Satisfying users’ demands can benefit the product from both competitiveness and user satisfaction. Feature requests have seen a rise in interest in the past few years and the amount of research has been growing. However, the diversity in the research topics suggests the need for their collective analysis to identify the challenges and opportunities so as to promote new advances in the future. In this work, following a defined process and a search protocol, we provide a systematic overview of the research area by searching and categorizing relevant studies. We select and analyze 140 primary studies using descriptive statistics and qualitative analysis methods. We classify the studies into different topics and group them from the perspective of requirements engineering activities. We investigate open tools as well as datasets for future research. In addition, we identify several key challenges and opportunities, such as: (1) ensuring the quality of feature requests, (2) improving their specification and validation, and (3) developing high-quality benchmarks for large language model-driven tasks. Our artifacts are publicly available at https://github.com/feifeiniu-se/FeatureRequestSurvey.
@@ 1. Introduction
Due to the considerable value of feature requests to software, the analysis and processing of feature requests in software has been widely studied in practice. Recent years have witnessed a surge research interest in feature requests, such as identifying feature requests from issues reports (Iacob and Harrison, 2013; Carreño and Winbladh, 2013; Deocadez et al., 2017; Dhinakaran et al., 2018), extracting software features from users’ requests (Bakar et al., 2016; Bakiu and Guzman, 2017; Johann et al., 2017; Licorish et al., 2017; Dalpiaz and Parente, 2019), feature requests prioritization (Kegel, 2023; de Lima and Marcacini, 2024; Salleh et al., 2024), and locating features in the source code (Palomba et al., 2017; Panichella et al., 2016; Ciurumelea et al., 2017; Kato and Goto, 2017). These have greatly promoted the application of feature requests in software engineering. Furthermore, to investigate the challenges and opportunities in software change request repositories, Cavalcanti et al. (Cavalcanti et al., 2014) conducted a systematic mapping study based on two dimensions: challenges and opportunities related to change request repositories. Tavakoli et al. (Tavakoli et al., 2018) surveyed the state-of-the-art (SOTA) tools and techniques employed in detecting informative aspects in mobile application reviews. Bakar et al. (Bakar et al., 2015) provided a systematic literature review of the SOTA approaches in feature extractions from natural language requirements for reuse in software product line engineering. Wang et al. (Wang et al., 2019) investigated the implicit feedback and explicit feedback employed for crowdsourced requirements engineering (RE) activities. Most existing secondary studies have focused on narrow aspects of feature requests, such as analyzing app store reviews for feature requests or addressing specific tasks like feature extraction. However, to the best of our knowledge, there is no comprehensive study that provides a holistic view of the entire research landscape on feature requests across different tasks and contexts. Furthermore, the field lacks an integrated discussion of overarching challenges and open research opportunities that span the full lifecycle of feature request management. Additionally, there is a lack of understanding regarding the frequency and location of studies published over the past few years.
@@
In this paper, we conduct a systematic literature review on feature requests to portray the landscape of research practices from a requirements engineering perspective and pinpoint research challenges and opportunities, following the guidelines of (Petersen et al., 2008, 2015). We propose four research questions and define a set of search strings to retrieve the initial relevant literature. Then snowballing is employed to retrieve the whole literature. Finally, a total of 140 literature studies are selected. In particular, the studies are scrutinized for feature requests analysis and processing for the software maintenance. We report on the demographics of the published articles to provide a comprehensive view. Furthermore, we also categorize the literature according to the research topic and provide a detailed analysis of each topic. Figure 1 presents an overview of the study topics in relation to requirements activities, encompassing 12 distinct research topics. We also expand each research topic with an explanation of the research progress, accompanied by a demonstration of the proposed approach. To facilitate future studies, we investigate the public tools and datasets for feature requests. Finally, we pinpoint the research gaps, which can help researchers understand the road ahead in investigating user requests.
@@ 2.2. Related Work
Cavalcanti et al. (Cavalcanti et al., 2014) performed a systematic mapping study in 2014 to characterize the research challenges and opportunities for software change request repositories. The result was a two-taxonomy based on two dimensions: challenges and opportunities related to change request repositories. However, the search strings of this paper mainly involve “change request” and “bug report”, etc., not “feature request”. Thus, some feature request-related research may be missed. Tavakoli et al. (Tavakoli et al., 2018) surveyed the SOTA tools and techniques used to detect valuable information concerning user needs and requirements from mobile application reviews. User reviews of mobile applications may contain bug reports, user experience, and user requirements, which are pivotal information relevant to app development and improvement. Therefore, they surveyed the mining tools that have been used to unleash the value of app reviews for mobile app development. They mainly focus on obtaining useful information from user reviews, which our study pays attention to the analysis and processing of user feature requests, which is downstream of their survey. Wang et al. (Wang et al., 2019) provided a mapping study of literature on crowdsourced user feedback employed for requirements engineering purposes. They provided an overview of the types of user feedback that have been employed for crowdsourced requirements engineering activities. They investigated both implicit feedback and explicit feedback. However, they only mapped the literature to requirements engineering activities without further investigation into the SOTA techniques for each topic. Bakar et al. (Bakar et al., 2015) provided a systematic literature review of the SOTA approaches in feature extractions from natural language requirements for reuse in software product line engineering. Dąbrowski et al. (Dąbrowski et al., 2022) conducted a systematic review of 182 studies on app review analysis for software engineering. They categorized existing work into feature request identification, bug report detection, sentiment analysis, and user experience evaluation, and summarized commonly used techniques such as machine learning, deep learning, and topic modeling. The study highlighted how app review analysis supports requirements, design, testing, and maintenance. However, their work remains a high-level categorization, lacking an in-depth methodological synthesis for each task. In contrast, our study discusses feature request analysis from the perspective of the software development lifecycle, providing a more comprehensive understanding of how feature requests are generated, processed, and utilized throughout different development stages. Dit et al. (Dit et al., 2013) conducted a comprehensive survey on feature location in source code, providing a detailed taxonomy that categorizes 89 studies according to multiple analytical dimensions, including static, dynamic, textual, historical, and hybrid approaches. The survey summarizes key techniques for locating the implementation of software features and identifies major research gaps, such as the lack of standardized benchmarks, inconsistent evaluation practices, and limited cross-language applicability. Bischoff et al. (Bischoff et al., 2019) created a panoramic view of the current literature to pinpoint gaps and supply insights in the integration of feature models. They sought to understand, characterize, and summarize the literature about the integration of feature models. Recently, Cai et al. (Cai et al., 2025) reviewed existing work on requirements elicitation from user-generated content, categorized studies by data sources, methods, and representations, and proposed a research framework to guide future work in this area.
Although several secondary studies have investigated concepts related to user feedback and software features, most have focused on adjacent topics or partial perspectives rather than offering a systematic synthesis of feature request research across the software development lifecycle. Existing work has examined areas such as change request repositories (Cavalcanti et al., 2014), requirement or information extraction from app reviews (Tavakoli et al., 2018; Cai et al., 2025), feature modeling (Bischoff et al., 2019), feature extraction in software product lines (Bakar et al., 2015), feature localization (Dit et al., 2013), and high-level mapping studies on crowdsourced feedback (Wang et al., 2019; Dąbrowski et al., 2022). However, these studies do not provide a unified view of how feature requests are managed throughout the software lifecycle.
In contrast, our study presents a lifecycle-oriented synthesis of feature request research. We categorize existing work according to key stages of the software development process and, within each stage, analyze the techniques and datasets employed. This perspective delivers a comprehensive and methodical understanding of how feature requests emerge, evolve, and contribute to continuous software evolution, offering a cohesive view that bridges fragmented prior research.
@@ 3.2.1. Search Strategy
As illustrated in Figure 3, we adopted a systematic search strategy combining both automated searching and snowballing to ensure comprehensive coverage of related studies while minimizing the risk of missing relevant work. To address our research questions, we first identified the following keywords related to the feature request topic:
Keywords related to feature request: feature request, user request, user requirement, enhancement report, enhancement request, just-in-time requirement
The same search keywords were applied across all digital libraries, with only minor syntactic adaptations to comply with each database’s specific query requirements. Two authors independently conducted the automated search, which helps avoid bias and ensures reliability. The search was performed across the digital libraries listed in Figure 3, covering all available publication years up to August 2025. After merging and removing duplicates, 17,381 distinct papers were obtained.
Figure 3. Study Search and Selection Process.
@@ 3.2.2. Inclusion/Exclusion Criteria
To identify the most relevant studies on feature request analysis and processing, we drew upon prior work (Giray et al., 2023; Zakari et al., 2020; Croft et al., 2022; Niu et al., 2025b; Hu et al., 2025) and established a set of inclusion (ICs) and exclusion criteria (ECs), summarized in Table 1. These criteria were applied to the titles, abstracts, and keywords of candidate papers to ensure relevance and quality.
@@
Through this rigorous screening process, we refined a comprehensive yet focused set of relevant studies for detailed analysis. Overall, our study pool consists of 101 selected papers.
Table 1. Inclusion and Exclusion Criteria.
@@
[…]
@@
[…]
@@
[…]
3.2.3. Snowballing
To ensure a comprehensive literature collection, we performed both forward and backward snowballing into our study selection process, following the guidelines outlined by Wohlin (Wohlin, 2014). This iterative process involved examining the references and citations of the selected studies in successive rounds. After completing the database search and study selection (resulting in 101 papers), we performed iterative forward and backward snowballing on the selected studies. As shown in Figure 3, a total of three rounds of snowballing were conducted, which retrieved 2,938, 1,741, and 259 papers in each round, respectively. These papers were screened using the same inclusion and exclusion criteria presented in Table 1, resulting in 37, 8, and 0 newly included papers in each iteration. Ultimately, 45 additional primary studies were retained, increasing the final number of included studies from 101 to 146. This process ensured that relevant studies that may have been missed during the automated search were thoroughly captured.
3.2.4. Quality Assessment
@@
Each criterion was assessed using a binary scoring scheme (satisfied = 1, not satisfied or unclear = 0). Four authors independently evaluated all 146 candidate studies. In cases of disagreement, the reviewers discussed the assessment results until consensus was reached. For each study, we calculated an overall quality score by summing the scores across all eight criteria, resulting in a maximum possible score of 8. Following common practices in secondary studies, we excluded papers that satisfied fewer than 70% of the quality criteria (i.e., a total score lower than 6). These studies typically lacked essential information on data sources, methodological details, or evaluation procedures, which would hinder reliable data extraction and synthesis. As a result of this quality assessment process, six papers were excluded for not meeting the established quality threshold, leaving 140 high-quality studies for detailed data extraction and further analysis. For transparency, the list of excluded papers and the corresponding unmet criteria are reported in our GitHub repository.
Table 2. Quality Assessment Checklist for Feature Request Studies.
@@
3.3. Data Extraction and Synthesis
Data Extraction
@@ 3.3. Data Extraction and Synthesis
Data Synthesis.
The goal of this survey is to aggregate evidence from primary studies and provide an overview of the state of the art on feature request analysis and processing. Following the systematic review guidelines (Keele et al., 2007), we synthesized the extracted data using a combination of descriptive (quantitative) synthesis and qualitative thematic synthesis, aligned with our four research questions. For RQ1 (Demographics), we performed descriptive synthesis on bibliographic and contextual attributes (e.g., publication year, venue type, and study type). We report the evolution of the literature over time and the distribution across venues using frequency counts. For RQ2 (Analysis and Processing Practices), we synthesized the extracted methodological information by grouping studies according to the main feature request tasks they target (e.g., classification, prioritization, summarization, management) and the approaches they employ. We summarize the prevalence of different practices (including preprocessing, feature extraction/representation, and evaluation setups) using structured comparison tables, and we narratively discuss commonalities and differences across studies. For RQ3 (Public Datasets and Tools), we first summarized availability-related information (e.g., whether datasets/tools are publicly accessible, their sources) in a structured catalog. We then conducted a qualitative categorization of the datasets and tools by their intended purposes and research themes, and reported results for each category to facilitate reuse. For RQ4 (Challenges and Opportunities), we conducted an inductive thematic synthesis. During data extraction, any text explicitly describing a limitation, difficulty, research gap, or future work was copied verbatim into the extraction form. Two reviewers independently performed open coding on these excerpts to identify recurring concepts and iteratively refined a shared codebook through discussion. Conceptually similar codes were merged into higher-level themes, which were manually reviewed to ensure consistent granularity and clear boundaries. We report the resulting themes together with representative evidence from primary studies and, where appropriate, the number of studies mentioning each theme. When divergent or conflicting observations emerged, we explicitly described them rather than forcing a single conclusion.
@@ 4. Demographics
A total of 140 research papers were included in the final study set, covering publications up to August 2025. As shown in Figure 4, the number of primary studies has increased significantly over time, showing a clear upward trend from 2011 to 2017, with the peak occurring in 2017 and 2019 (16 studies). Activity remained relatively high through 2020, followed by a decline likely due to shifting research priorities. A notable resurgence occurred in 2024 (11 studies), coinciding with the rise of LLMs, which introduced new research opportunities. Overall, the data reveals cyclical research activity with notable peaks and dips over the years. Figure 5 presents the distribution of the selected papers across journals and conferences, highlighting the most frequently appearing publication venues. A total of 88 papers were published in conferences, 46 in journals, and 6 as preprints. Most studies were published in top-tier software engineering outlets, particularly in the field of requirements engineering. High-frequency venues include International Conference on Requirements Engineering (RE Conf., 15 papers), International Conference on Automated Software Engineering (ASE Conf., 8 papers), International Working Conference on Requirement Engineering: Foundation for Software Quality (REFSQ, 7papers), International Conference on Software Engineering (ICSE, 7 papers), Empirical Software Engineering (ESE, 7 papers), Journal of Systems and Software (JSS, 6 papers), Transactions on Software Engineering (TSE, 4 papers), International Conference on Software Maintenance and Evolution (ICSME, 4 papers), Requirements Engineering Journal (RE Jour., 4 papers), and International Conference on Evaluation and Assessment in Software Engineering (EASE, 4 papers). These results confirm that the topic has received strong attention from the software engineering community.
@@ 5. Analyzing and Processing Feature Requests
Throughout the lifecycle of feature requests, twelve research topics have been identified in the literature on feature request analysis and processing: feature request identification, quality review, approval prediction, feature extraction, intent recognition, summarization, classification, prioritization, traceability, developer assignment, localization, and code recommendation, covering the entire process.